# OODEval: Evaluating Large Language Models on Object-Oriented Design

### v1

@@ 5.2. Grammatical Correctness Evaluation Metric
5.3. Prompt Design
In this section, we detail the prompting strategy employed to enable LLMs to address each OOD task within the OODEval benchmark. To promote equitable evaluation across models, we adopted methodologies akin to those outlined in (Du et al., 2024; Ding et al., 2025), thereby developing a standardized prompt template compatible with both instruction-fine-tuned and base LLMs. The template comprises the following elements: First, a system role prompt defines the model’s assumed persona and delineates the overarching task objectives. This is followed by task-specific instructions, which encompass a description of the requirements and detailed guidance for the given task. The model is directed to produce PlantUML code derived from the requirement description, delimited by @startuml and @enduml markers to streamline code extraction in post-processing. Furthermore, the instructions mandate that the generated PlantUML code explicitly incorporate class names, attributes, methods, and inter-class relationships. To enhance compatibility with base models and facilitate more accurate next-token prediction, we appended the marker ”##Response” at the conclusion of the prompt.
Prompt for Object-Oriented Design ## System Message
you are a professional UML class diagram design expert, which can generate the corresponding PlantUML class diagram design based on system requirements.
## Instruction
System requirement is as follows :
$\{user\_requirement\}$
please generate plantuml code based on system requirement. You should follow the instructions below:
(1) Generate standard PlantUML class diagram code directly, start with @startuml and end with @enduml tags. Do not generate any analysis, explanations, or irrelevant content. (2) Class names, attributes, and method names should use meaningful english names from system requirement. (3) Reasonably use inheritance, implementation, dependence, association, aggregation, and composition relationships to design the class diagram. ## Response
@@ 5.4. Implementation Details
For large-scale online LLMs, such as GPT-4o and GPT-3.5-turbo, we accessed the models via their respective API interfaces. For open-source local LLMs, we downloaded and executed their released versions directly from the official repositories, following the provided documentation. To ensure consistency, the maximum context window length was uniformly set to 4096 tokens across all models, corresponding to the smallest supported window among the evaluated LLMs. All experiments were conducted on a server equipped with 8 NVIDIA L40S GPUs, each with 46 GB of memory. For each task, we generated five samples of design solutions, yielding a total of 7,250 outputs (29 LLMs × 50 tasks × 5 samples). Following prior studies (Chen et al., 2021; Du et al., 2024; Cao et al., 2024), we adopted a temperature of 0.2, default values for top-$p$ and top-$k$, and a maximum output token length of 2048 to produce complete responses without truncation. The datasets, source code, and experimental results of this study are publicly available in a reproducibility package on Zenodo (OODEval, 2026).
@@ 6.1. RQ1: Overall Correctness
Figure 3. Performance Comparison of Pass@1 and CLUE Metrics.
@@ 6.2. RQ2: Comparison With Human
First, comparing the average performance of humans to that of LLMs reveals that LLMs significantly underperform the human average across all CLUE metrics. The most pronounced disparities occur in class method generation (clue-method, exceeding a 50% gap) and class relationship identification (clue-relation, exceeding a 30% gap). Collectively, these results indicate that the average LLMs remain substantially inferior to the average human proficiency in OOD tasks.
Second, when benchmarking the human average against the best LLM performance, notable findings emerge. The top model for attribute generation (Llama3.1-70B-Instruct) marginally surpasses the human average, while the leading model for relationship generation (Gemma3-27B-Instruct) approximates it closely. In contrast, the strongest performer across the overall clue, clue-class, and clue-method metrics (Qwen3-Coder-30B) falls slightly below the human average. These deviations are minimal, not exceeding 5 percentage points. Thus, the premier LLMs achieve near parity with average human performance.
Finally, contrasting the peak human performance with the best LLMs highlights a clear disparity. The highest-performing human participants accomplish OOD tasks with near-perfect accuracy, attaining scores above 99% on all CLUE metrics. Evidently, even the most advanced LLMs currently exhibit a considerable gap relative to the best human benchmark.
@@ 6.3. RQ3: Model Dimension Analysis
Finally, we compared reasoning LLMs(RLMs) with non-reasoning LLMs(non-RLMs) using four model pairs, as depicted in Figure 6(d). Qwen3 (released April 2025) is reasoning-enhanced, while Gemma3 (released March 2025) is not. We paired Qwen3 and Gemma3 models of comparable parameter scales for comparison. Additionally, DeepSeek-R1 and DeepSeek-V3 were included as another control pair. The results reveal that Gemma3 outperforms Qwen3, and DeepSeek-V3 trails DeepSeek-R1 by just 1 percentage point. Thus, reasoning models show no clear advantage and may underperform compared to their non-reasoning counterparts. However, since we did not ensure identical underlying model architectures and model parameter scale across pairs, these findings should be regarded as preliminary observations.

### v2

@@ 5.1. Studied LLMs
5.2. Prompt Design
In this section, we describe the prompt design used to enable LLMs to perform each OOD task in the OODEval benchmark. To ensure fair comparison across models, we follow prior benchmarking practices (Du et al., 2024; Ding et al., 2025) and adopt a unified prompt template applicable to both instruction-tuned and base LLMs. The template consists of a system role prompt that specifies the model’s assumed role and overall task goal, followed by task-specific instructions detailing the requirement description and expected output. Models are instructed to generate PlantUML code enclosed by @startuml and @enduml markers to facilitate automatic extraction, with explicit inclusion of class names, attributes, methods, and inter-class relationships.
Prompt for Object-Oriented Design ## System Message you are a professional UML class diagram design expert, which can generate the corresponding PlantUML class diagram design based on system requirements. ## Instruction System requirement is as follows : $\{user\_requirement\}$ please generate plantuml code based on system requirement. You should follow the instructions below: (1) Generate standard PlantUML class diagram code directly, start with @startuml and end with @enduml tags. Do not generate any analysis, explanations, or irrelevant content. (2) Class names, attributes, and method names should use meaningful english names from system requirement. (3) Reasonably use inheritance, implementation, dependence, association, aggregation, and composition relationships to design the class diagram. ## Response
@@ 5.4. Implementation Details
For large-scale online LLMs, such as GPT-4o and GPT-3.5-turbo, we accessed the models via their respective API interfaces. For open-source local LLMs, we downloaded and executed their released versions directly from the official repositories, following the provided documentation. To ensure consistency, the maximum context window length was uniformly set to 4096 tokens across all models, corresponding to the smallest supported window among the evaluated LLMs. All experiments were conducted on a server equipped with 8 NVIDIA L40S GPUs, each with 46 GB of memory. For each task, we generated five samples of design solutions, yielding a total of 7,250 outputs (29 LLMs × 50 tasks × 5 samples). Following prior studies (Chen et al., 2021; Du et al., 2024; Cao et al., 2024), we adopted a temperature of 0.2, default values for top-$p$ and top-$k$, and a maximum output token length of 2048 to produce complete responses without truncation.
@@ 6.1. RQ1: Overall Correctness
Figure 3. Performance Comparison of Pass@1 and CLUE Metrics.
@@ 6.2. RQ2: Comparison With Human
First, comparing the average performance of humans to that of LLMs reveals that LLMs significantly underperform the human average across all CLUE metrics. The most pronounced disparities occur in class method generation (clue-method, exceeding a 50% gap) and class relationship identification (clue-relation, exceeding a 30% gap). Collectively, these results indicate that the average LLMs remain substantially inferior to the average human proficiency in OOD tasks. Second, when benchmarking the human average against the best LLM performance, notable findings emerge. The top model for attribute generation (Llama3.1-70B-Instruct) marginally surpasses the human average, while the leading model for relationship generation (Gemma3-27B-Instruct) approximates it closely. In contrast, the strongest performer across the overall clue, clue-class, and clue-method metrics (Qwen3-Coder-30B) falls slightly below the human average. These deviations are minimal, not exceeding 5 percentage points. Thus, the premier LLMs achieve near parity with average human performance. Finally, contrasting the peak human performance with the best LLMs highlights a clear disparity. The highest-performing human participants accomplish OOD tasks with near-perfect accuracy, attaining scores above 99% on all CLUE metrics. Evidently, even the most advanced LLMs currently exhibit a considerable gap relative to the best human benchmark.
@@ 6.3. RQ3: Model Dimension Analysis
Finally, we compared reasoning LLMs(RLMs) with non-reasoning LLMs(non-RLMs) using four model pairs, as depicted in Figure 6(d). Qwen3 (released April 2025) is reasoning-enhanced, while Gemma3 (released March 2025) is not. We paired Qwen3 and Gemma3 models of comparable parameter scales for comparison. Additionally, DeepSeek-R1 and DeepSeek-V3 were included as another control pair. The results reveal that Gemma3 outperforms Qwen3, and DeepSeek-V3 trails DeepSeek-R1 by just 1 percentage point. Thus, reasoning models show no clear advantage and may underperform compared to their non-reasoning counterparts. Since we did not ensure identical underlying model architectures and model parameter scale across pairs, this finding should be regarded as preliminary observations.
@@ 9. Conclusion
Acknowledgments
This work was supported by the Major Program of the National Natural Science Foundation of China (Grant Nos. 62192733, 62192730) and the National Natural Science Foundation of China (Grant No. 62372219).